# Dynamics and lifetime of geometric excitations in moiré systems

### v1

@@ VI Experimental discussions based on tTMD systems
Such interaction is the Zhang-Das-Sarma (ZDS) potential, commonly used to describe finite-thickness effects in QH systems [99, 100]. This interaction has smaller short-range components than the standard Coulomb interaction, making it effective in reducing the GM energy. Here, $\lambda\cdot a_{M}$ represents the “thickness” of the wavefunctions in tMoTe2, accounting for the spread of electronic wavefunctions in the out-of-plane direction. As shown in Fig. 4c, a slight increase in $\lambda$ shifts the GM peak downward from the continuum boundary and enhances its magnitude compared to the pure Coulomb interaction as confirmed by Fig. 4b. Experimentally, this increase in the effective thickness of the tMoTe2 sample can be achieved by inserting a thin dielectric medium between the twisted layers or using twisted multilayer MoTe2. However, it is crucial to maintain the interlayer distance within an optimal range, as an excessive increase in $\lambda$ may induce phase transitions that destabilize the FCI phase, which requires further investigation and will be left for future studies.

### v2

@@ VI Experimental discussions based on tTMD systems
Figure 6: Spectral functions and spectra of modified Coulomb interaction $V_{\text{ZDS}}$ with effective thickness $\lambda=0.2$ in tMoTe2. For the fermionic system (a, b) with the filling $\nu=-2/3$, the graviton mode (GM) peak is enhanced by approximately 50% compared to the pure Coulomb case (shown in the supplementary materials). However, the peak strength still decreases as the system size increases. In panel b, the GM position (red dot) lies slightly below the continuum. In contrast, the bosonic system (c, d) with $\nu=-1/2$ exhibits a more robust response. The spectral peak remains strong and even increases slightly with system size. Moreover, the GM position is fully tunable to fall within the magnetoroton modes and becomes well-separated from the continuum. The hole number in b and d is $N_{h}=10$, where $(K_{1},K_{2})$ gives the position in the first Brillouin zone with $N_{1}$ as the number of sites along one of the primitive lattice vectors.
@@
Such interaction is the Zhang-Das-Sarma (ZDS) potential, commonly used to describe finite-thickness effects in QH systems [100, 101]. This interaction has smaller short-range components than the standard Coulomb interaction, making it effective in reducing the GM energy. Here, $\lambda\cdot a_{M}$ represents the “thickness” of the wavefunctions in tMoTe2, accounting for the spread of electronic wavefunctions in the out-of-plane direction. As shown in Fig. 6b, a slight increase in $\lambda$ shifts the GM peak downward from the continuum boundary and enhances its magnitude compared to the pure Coulomb interaction as confirmed by Fig. 6a  [^5] The results for pure Coulomb and more $\lambda$ values are shown in the supplementary material. Experimentally, this increase in the effective thickness of the tMoTe2 sample can be achieved by inserting a thin dielectric medium between the twisted layers or using twisted multilayer MoTe2.
One concern arising from Fig. 6a is that the GM peak continues to decrease with increasing system size because the GM is not significantly away from the continuum, making it uncertain whether the proposed tuning strategy can sustain the GM peak in the thermodynamic limit. Drawing from insights gained in LLs, we propose a more robust approach for probing GMs in moiré Chern bands by utilizing bosonic systems. In bosonic FQH phases, the magnetoroton modes tend to be flatter, and the GM modes generally have lower energy, increasing the likelihood of achieving full separation from the continuum [103, 79, 9].
Experimentally, bosonic FQH phases can be realized using ultracold atoms in optical lattices [104, 105, 106, 107] or interacting photons [108]. If the quantum geometric properties of the moiré lattice can be encoded into an experimental system featuring bosonic FCI phases with tunable interactions, this approach will provide significant advantages. Fig. 6c presents the spectral function for the ZDS interaction with $\lambda=0.2$ in tMoTe2, demonstrating that the GM peak is significantly enhanced compared to the fermionic case, with minimal decay as the system size increases. Furthermore, Fig. 6d confirms that the GM is fully gapped from the continuum and falls within the magnetoroton mode, further supporting the robustness of this approach. Although the realization of bosonic FQH phases is in an early stage, these results provide compelling evidence for the feasibility of probing GMs in bosonic moiré Chern bands as experimental techniques continue to advance. However, regardless of particle statistics, it is essential to maintain the effective interlayer distance $\lambda$ within an optimal range. A significant increase in $\lambda$ may induce phase transitions that destabilize the FCI phase, which requires further investigation and will be left for future studies.
VII Conclusion and outlook
By analyzing the spectral functions of the chiral graviton operators in moiré lattice Chern bands, we find that GMs become significantly weaker than in LLs. Similar behaviors are also seen with GMs generated with SMA in the long wavelength limit. Specifically, the GM lifetime in lattice Chern bands decreases rapidly with an increase in the density of states, a phenomenon further exacerbated when the GM resides within the excitation continuum. This behavior is well captured by a minimal theoretical model, which reveals the mismatch between the symmetries of the ground state and the excited states in generic lattice Chern bands. Moreover, we propose that probing the graviton mode could be achieved by further suppressing the short-range component of the potential to position the GM below the continuum or using a bosonic system that can realize the FCI phases with moiré potential. Additionally, we suggest a potential tuning strategy to enhance the GM resonance peak in experimental settings.
Despite this progress, many questions remain regarding the GM and other neutral excitations in moiré materials. One key question is whether the graviton mode can be shifted below the continuum in experiments by tuning the effective interactions without inducing a phase transition. Another promising direction is to explore other neutral collective modes in Chern bands, such as finite-momentum GMP modes, gravitino modes, higher-spin modes, and the multiple graviton modes that emerge at different filling fractions in fractional quantum Hall systems. One can also look into the counterparts of higher LLs’ neutral excitations in the moiré systems. Furthermore, the extremely short lifetime of GMs may indicate a breakdown of the single-mode approximation in the long-wavelength limit of lattice Chern bands. This suggests the possible emergence of novel low-lying neutral excitations in FCI phases, which requires deeper investigation and understanding.